OpenStudy (anonymous):

Keisha's penny bank is 7/10 full. After she removes 200 pennies, it is 1/2 full. How many pennies can Keisha's bank hold?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what is \[\frac{7}{10}-\frac{1}{2}\]?

OpenStudy (anonymous):

1/5

OpenStudy (misty1212):

ok good so 200 is one fifth of the total the bank can hold

OpenStudy (anonymous):

if 200 is half then what is 200*2

OpenStudy (anonymous):

or 200+200

OpenStudy (misty1212):

the total should therefore be \[200\times 5\]

OpenStudy (misty1212):

@sweetness221 200 is NOT half

OpenStudy (anonymous):

oh sry i understood the question wrong

OpenStudy (anonymous):

thank you @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]